OpenStudy (anonymous):

ms. burke invested $25,000 in two accounts, one yielding 4% interest and the other yielding 8%. if she received a total of $1,800 in interest at the end of the year, how much did she invest in each account?

OpenStudy (whpalmer4):

You've got two funds, fund A and fund B. Fund A pays 4%, and B pays 8%. Total investment = $25,000 = A + B A*4% + B*8% = $1800 A + B = 25,000 0.04A + 0.08B = 1800 (or 4A + 8B = 180000) Solve for A and B via substitution or elimination.

OpenStudy (whpalmer4):

If you want to approach this from a back of the envelope POV, investing the entire $25,000 at 8% would yield $2000, and investing the entire $25,000 at 4% would yield only $1000, so we know that most of the money goes into the B fund. Because our target figure of $1800 is 4/5 of the way between $1000 and $2000, we know that 4/5 of the money must go into the higher yielding account.